OpenStudy (anonymous):

Use differentials to estimate the value indicated: 1001^(1/3)

OpenStudy (anonymous):

So f(1001)~f(1000)+f'(1000)*(?)

OpenStudy (anonymous):

function is \[f(x)=\sqrt[3]{1000+x}\] linearize that one, replace \(x\) by 1

OpenStudy (anonymous):

The question mark is where I'm stuck. What do I multiple derivative by?

OpenStudy (anonymous):

equation of line tangent to the curve at the point \((1000,10)\) is \[y-10=m(x-1000)\] and you find \(m\) by taking the derivative and replacing \(x\) by 1000

OpenStudy (anonymous):

or better yet, use the function i wrote and the point \((0,10)\)